OpenStudy (anonymous):

Find dy/dx of y^2 + xy = 1 Help please!! :(

OpenStudy (abb0t):

2y+xdy/dx+y = 0

OpenStudy (anonymous):

dy/dx=y^2+xy=1 2y(dy/dx)+xy(dy/dx)=0 2y+xy(dy/dx)=0 dy/dx=0

OpenStudy (anonymous):



OpenStudy (abb0t):

2ydy/dx*

OpenStudy (anonymous):